OpenStudy (aaronandyson):

Why infinity minus infinity not equal to 0?

OpenStudy (marcelie):

i think because the limit is too large

OpenStudy (aaronandyson):

Suppose, \[1 -\frac{ 1 }{ 2 } + \frac{ 1 }{ 3 } - \frac{ 1 }{ 4 } + \frac{ 1 }{ 5 } - \frac{ 1 }{ 6 }...\]

myininaya (myininaya):

infinity's can differ (like they aren't all the same "size") for example one infinity could be "larger" than another infinity \[\lim_{x \rightarrow \infty}x =\infty \\ \lim_{x \rightarrow \infty}x^2=\infty \\ \text{ and } \\ \lim_{x \rightarrow \infty} (x-x^2)=-\infty \]

OpenStudy (aaronandyson):

what if, \[\infty-\infty+1 =0+1 = 1?\]

OpenStudy (l):

Ohhh, that's a very good example.\[\sum_{r=1}^{\infty} \frac{1}{2r-1} - \sum_{r=1}^{\infty} \frac{1}{2r} = \ln (2)\]

OpenStudy (skullpatrol):

Infinity is not a number.

OpenStudy (aaronandyson):

Then what is it?

OpenStudy (skullpatrol):

Like all words it depends on your definition to give it meaning.

myininaya (myininaya):

i like the saying that all infinities aren't created equal

myininaya (myininaya):

I can't remember where I heard that from

myininaya (myininaya):

like one function could get infinitely bigger faster than another function like in my example

Nnesha (nnesha):

\[\huge\rm \infty \cancel{=} \infty\] @skullpatrol xP

OpenStudy (skullpatrol):

$$\infty - \infty = ?$$

myininaya (myininaya):